# TimeBridge: Non-Stationarity Matters for Long-term Time Series Forecasting

### v3

@@ E.1 Main Experiments
Furthermore, on the large Climate dataset (Godahewa et al., 2021), which contains 1,763 channels, the prediction task becomes particularly challenging due to the more complex inter-channel dependencies and higher levels of noise in the data. Despite these challenges, TimeBridge maintains its superior performance over competing models.
@@
TimeMixer PatchTST Crossformer FEDformer ModernTCN MICN TimesNet DLinear
@@
ETTm1 96 0.284 0.337 0.300 0.353 0.291 0.347 0.293 0.345 0.293 0.346 0.310 0.361 0.326 0.390 0.292 0.346 0.314 0.360 0.338 0.375 0.299 0.343
192 0.317 0.367 0.345 0.382 0.325 0.372 0.335 0.372 0.333 0.370 0.363 0.402 0.365 0.415 0.332 0.368 0.359 0.387 0.371 0.387 0.335 0.365
336 0.361 0.394 0.374 0.398 0.359 0.390 0.368 0.386 0.369 0.392 0.408 0.430 0.392 0.425 0.365 0.391 0.398 0.413 0.410 0.411 0.369 0.386
720 0.413 0.418 0.429 0.430 0.418 0.423 0.426 0.417 0.416 0.420 0.600 0.547 0.446 0.458 0.416 0.417 0.459 0.464 0.478 0.450 0.425 0.421
Avg. 0.344 0.379 0.362 0.391 0.348 0.383 0.355 0.380 0.353 0.382 0.420 0.435 0.382 0.422 0.351 0.381 0.383 0.406 0.400 0.406 0.357 0.379
ETTm2 96 0.157 0.243 0.175 0.266 0.169 0.258 0.165 0.256 0.166 0.256 0.263 0.359 0.180 0.271 0.166 0.256 0.178 0.273 0.187 0.267 0.167 0.260
192 0.217 0.285 0.242 0.312 0.229 0.299 0.225 0.298 0.223 0.296 0.345 0.400 0.252 0.318 0.222 0.293 0.245 0.316 0.249 0.309 0.224 0.303
336 0.269 0.321 0.282 0.340 0.280 0.333 0.277 0.332 0.274 0.329 0.469 0.496 0.324 0.364 0.272 0.324 0.295 0.350 0.321 0.351 0.281 0.342
720 0.348 0.378 0.378 0.398 0.349 0.384 0.360 0.387 0.362 0.385 0.996 0.750 0.410 0.420 0.351 0.381 0.389 0.406 0.497 0.403 0.397 0.421
Avg. 0.246 0.310 0.269 0.329 0.257 0.319 0.257 0.318 0.256 0.317 0.518 0.501 0.292 0.343 0.253 0.314 0.277 0.336 0.291 0.333 0.267 0.332
ETTh1 96 0.350 0.389 0.386 0.405 0.369 0.396 0.372 0.401 0.370 0.400 0.386 0.426 0.376 0.415 0.368 0.394 0.396 0.427 0.384 0.402 0.375 0.399
192 0.388 0.414 0.424 0.440 0.410 0.417 0.413 0.430 0.413 0.429 0.413 0.442 0.423 0.446 0.405 0.413 0.430 0.453 0.557 0.436 0.405 0.416
336 0.408 0.430 0.449 0.460 0.391 0.414 0.438 0.450 0.422 0.440 0.440 0.461 0.444 0.462 0.391 0.412 0.433 0.458 0.491 0.469 0.439 0.443
720 0.443 0.463 0.495 0.487 0.447 0.464 0.486 0.484 0.447 0.468 0.519 0.524 0.469 0.492 0.450 0.461 0.474 0.508 0.521 0.500 0.472 0.490
Avg. 0.397 0.424 0.439 0.448 0.404 0.423 0.427 0.441 0.413 0.434 0.440 0.463 0.428 0.454 0.404 0.420 0.433 0.462 0.458 0.450 0.423 0.437
ETTh2 96 0.271 0.331 0.297 0.348 0.272 0.334 0.281 0.351 0.274 0.337 0.611 0.557 0.332 0.374 0.263 0.332 0.289 0.357 0.340 0.374 0.289 0.353
192 0.335 0.370 0.371 0.403 0.325 0.369 0.349 0.387 0.314 0.382 0.703 0.624 0.407 0.446 0.320 0.374 0.409 0.438 0.402 0.414 0.383 0.418
336 0.371 0.402 0.404 0.428 0.319 0.373 0.366 0.413 0.329 0.384 0.827 0.675 0.400 0.447 0.313 0.376 0.417 0.452 0.452 0.452 0.448 0.465
720 0.387 0.425 0.424 0.444 0.395 0.433 0.401 0.436 0.379 0.422 1.094 0.775 0.412 0.469 0.392 0.433 0.426 0.473 0.462 0.468 0.605 0.551
Avg. 0.341 0.382 0.374 0.406 0.328 0.377 0.349 0.397 0.324 0.381 0.809 0.658 0.388 0.434 0.322 0.379 0.385 0.430 0.414 0.427 0.431 0.447
Weather 96 0.144 0.184 0.159 0.208 0.146 0.198 0.147 0.198 0.149 0.198 0.146 0.212 0.217 0.296 0.149 0.200 0.161 0.226 0.172 0.220 0.152 0.237
192 0.186 0.225 0.200 0.248 0.191 0.239 0.192 0.243 0.194 0.241 0.195 0.261 0.275 0.329 0.196 0.245 0.220 0.283 0.219 0.261 0.220 0.282
336 0.237 0.267 0.253 0.289 0.241 0.280 0.247 0.284 0.245 0.282 0.252 0.311 0.339 0.377 0.238 0.277 0.275 0.328 0.280 0.306 0.265 0.319
720 0.307 0.320 0.321 0.338 0.310 0.331 0.318 0.330 0.314 0.334 0.318 0.363 0.389 0.409 0.314 0.334 0.311 0.356 0.365 0.359 0.323 0.362
Avg. 0.219 0.249 0.233 0.271 0.222 0.262 0.226 0.264 0.226 0.264 0.228 0.287 0.305 0.287 0.224 0.264 0.242 0.298 0.259 0.287 0.240 0.300
Electricity 96 0.120 0.214 0.138 0.237 0.132 0.234 0.153 0.256 0.129 0.222 0.135 0.237 0.183 0.297 0.129 0.226 0.159 0.267 0.168 0.272 0.140 0.237
192 0.142 0.237 0.157 0.256 0.148 0.248 0.168 0.269 0.147 0.240 0.160 0.262 0.195 0.308 0.143 0.239 0.168 0.279 0.184 0.289 0.152 0.249
336 0.156 0.252 0.167 0.264 0.165 0.266 0.189 0.291 0.163 0.259 0.182 0.282 0.212 0.313 0.161 0.259 0.196 0.308 0.198 0.300 0.169 0.267
720 0.179 0.278 0.194 0.286 0.197 0.296 0.228 0.320 0.197 0.290 0.246 0.337 0.231 0.343 0.191 0.286 0.203 0.312 0.220 0.320 0.203 0.301
Avg. 0.149 0.245 0.164 0.261 0.161 0.261 0.185 0.284 0.159 0.253 0.181 0.279 0.205 0.315 0.156 0.253 0.182 0.292 0.192 0.295 0.166 0.264
Traffic 96 0.340 0.240 0.363 0.265 0.355 0.261 0.369 0.257 0.360 0.249 0.512 0.282 0.562 0.349 0.368 0.253 0.508 0.301 0.593 0.321 0.410 0.282
192 0.343 0.250 0.385 0.273 0.380 0.271 0.400 0.272 0.379 0.256 0.501 0.273 0.562 0.346 0.379 0.261 0.536 0.315 0.617 0.336 0.423 0.287
336 0.363 0.257 0.396 0.277 0.393 0.281 0.407 0.272 0.392 0.264 0.507 0.279 0.570 0.323 0.397 0.270 0.525 0.310 0.629 0.336 0.436 0.296
720 0.393 0.271 0.445 0.312 0.434 0.300 0.461 0.316 0.432 0.286 0.571 0.301 0.596 0.368 0.440 0.296 0.571 0.323 0.640 0.350 0.466 0.315
Avg. 0.360 0.255 0.397 0.282 0.391 0.278 0.409 0.279 0.391 0.264 0.523 0.284 0.573 0.347 0.396 0.270 0.535 0.312 0.620 0.336 0.434 0.295
Solar 96 0.161 0.224 0.188 0.242 0.165 0.238 0.179 0.232 0.178 0.229 0.166 0.230 0.201 0.304 0.202 0.263 0.188 0.252 0.219 0.314 0.216 0.287
192 0.177 0.237 0.193 0.258 0.184 0.254 0.201 0.259 0.189 0.246 0.186 0.237 0.237 0.337 0.223 0.279 0.215 0.280 0.231 0.322 0.244 0.305
336 0.188 0.244 0.195 0.259 0.191 0.263 0.190 0.256 0.198 0.249 0.203 0.243 0.254 0.362 0.241 0.292 0.222 0.267 0.246 0.337 0.263 0.319
720 0.197 0.252 0.223 0.281 0.199 0.262 0.203 0.261 0.209 0.256 0.210 0.256 0.280 0.397 0.247 0.292 0.226 0.264 0.280 0.363 0.264 0.324
Avg. 0.181 0.239 0.200 0.260 0.185 0.254 0.193 0.252 0.194 0.245 0.191 0.242 0.243 0.350 0.228 0.282 0.213 0.266 0.244 0.334 0.247 0.309
Climate 96 1.040 0.492 1.044 0.494 1.042 0.498 1.341 0.578 1.354 0.519 1.083 0.508 1.290 0.620 1.100 0.514 1.150 0.574 1.274 0.524 1.108 0.536
192 1.049 0.494 1.055 0.496 1.051 0.494 1.447 0.622 1.388 0.572 1.115 0.532 1.447 0.693 1.125 0.523 1.159 0.568 1.274 0.541 1.122 0.536
336 1.066 0.495 1.069 0.495 1.072 0.496 1.430 0.610 1.410 0.581 1.246 0.575 1.293 0.625 1.153 0.530 1.173 0.571 1.285 0.584 1.146 0.545
720 1.071 0.496 1.073 0.497 1.075 0.497 1.438 0.615 1.395 0.575 1.236 0.584 1.480 0.710 1.243 0.599 1.183 0.568 1.347 0.626 1.158 0.545
Avg. 1.057 0.494 1.060 0.496 1.060 0.496 1.414 0.606 1.387 0.562 1.170 0.550 1.378 0.662 1.155 0.542 1.166 0.570 1.295 0.569 1.134 0.541
@@ E.2 Ablation Studies
Appendix F Statistical Analysis
@@
Climate $1.057\pm 0.012$ $0.494\pm 0.004$ $1.060\pm 0.011$ $0.496\pm 0.006$ $99\%$
Table 15: Standard deviation and statistical tests for TimeBridge and second-best method (DeformableTST) on ETT, Weather, Electricity, Traffic, Solar and Climate datasets.
@@
Dataset ARR AVol MDD ASR CR IR ARR AVol MDD ASR CR IR Interval

### v4

@@ F.1 Main Experiments
TimeMixer PatchTST Crossformer Leddam ModernTCN MICN TimesNet DLinear
@@
ETTm1 96 0.284 0.337 0.300 0.353 0.291 0.347 0.293 0.345 0.293 0.346 0.310 0.361 0.294 0.347 0.292 0.346 0.314 0.360 0.338 0.375 0.299 0.343
192 0.317 0.367 0.345 0.382 0.325 0.372 0.335 0.372 0.333 0.370 0.363 0.402 0.334 0.370 0.332 0.368 0.359 0.387 0.371 0.387 0.335 0.365
336 0.361 0.394 0.374 0.398 0.359 0.390 0.368 0.386 0.369 0.369 0.389 0.430 0.392 0.425 0.365 0.391 0.398 0.413 0.410 0.411 0.369 0.386
720 0.413 0.418 0.429 0.430 0.418 0.423 0.426 0.417 0.416 0.420 0.600 0.547 0.421 0.419 0.416 0.417 0.459 0.464 0.478 0.450 0.425 0.421
Avg. 0.344 0.379 0.362 0.391 0.348 0.383 0.355 0.380 0.353 0.382 0.420 0.435 0.354 0.381 0.351 0.381 0.383 0.406 0.400 0.406 0.357 0.379
ETTm2 96 0.157 0.243 0.175 0.266 0.169 0.258 0.165 0.256 0.166 0.256 0.263 0.359 0.174 0.260 0.166 0.256 0.178 0.273 0.187 0.267 0.167 0.260
192 0.217 0.285 0.242 0.312 0.229 0.299 0.225 0.298 0.223 0.296 0.345 0.400 0.231 0.301 0.222 0.293 0.245 0.316 0.249 0.309 0.224 0.303
336 0.269 0.321 0.282 0.340 0.280 0.333 0.277 0.332 0.274 0.329 0.469 0.496 0.288 0.336 0.272 0.324 0.295 0.350 0.321 0.351 0.281 0.342
720 0.348 0.378 0.378 0.398 0.349 0.384 0.360 0.387 0.362 0.385 0.996 0.750 0.368 0.386 0.351 0.381 0.389 0.406 0.497 0.403 0.397 0.421
Avg. 0.246 0.310 0.269 0.329 0.257 0.319 0.257 0.318 0.256 0.317 0.518 0.501 0.265 0.320 0.253 0.314 0.277 0.336 0.291 0.333 0.267 0.332
ETTh1 96 0.350 0.389 0.386 0.405 0.369 0.396 0.372 0.401 0.370 0.400 0.386 0.426 0.377 0.394 0.368 0.394 0.396 0.427 0.384 0.402 0.375 0.399
192 0.388 0.414 0.424 0.440 0.410 0.417 0.413 0.430 0.413 0.429 0.413 0.442 0.408 0.427 0.405 0.413 0.430 0.453 0.557 0.436 0.405 0.416
336 0.408 0.430 0.449 0.460 0.391 0.414 0.438 0.450 0.422 0.440 0.440 0.461 0.424 0.437 0.391 0.412 0.433 0.458 0.491 0.469 0.439 0.443
720 0.443 0.463 0.495 0.487 0.447 0.464 0.486 0.484 0.447 0.468 0.519 0.524 0.451 0.465 0.450 0.461 0.474 0.508 0.521 0.500 0.472 0.490
Avg. 0.397 0.424 0.439 0.448 0.404 0.423 0.427 0.441 0.413 0.434 0.440 0.463 0.415 0.430 0.404 0.420 0.433 0.462 0.458 0.450 0.423 0.437
ETTh2 96 0.271 0.331 0.297 0.348 0.272 0.334 0.281 0.351 0.274 0.337 0.611 0.557 0.283 0.345 0.263 0.332 0.289 0.357 0.340 0.374 0.289 0.353
192 0.335 0.370 0.371 0.403 0.325 0.369 0.349 0.387 0.314 0.382 0.703 0.624 0.339 0.381 0.320 0.374 0.409 0.438 0.402 0.414 0.383 0.418
336 0.371 0.402 0.404 0.428 0.319 0.373 0.366 0.413 0.329 0.384 0.827 0.675 0.366 0.405 0.313 0.376 0.417 0.452 0.452 0.452 0.448 0.465
720 0.387 0.425 0.424 0.444 0.395 0.433 0.401 0.436 0.379 0.422 1.094 0.775 0.395 0.436 0.392 0.433 0.426 0.473 0.462 0.468 0.605 0.551
Avg. 0.341 0.382 0.374 0.406 0.328 0.377 0.349 0.397 0.324 0.381 0.809 0.658 0.345 0.391 0.322 0.379 0.385 0.430 0.414 0.427 0.431 0.447
Weather 96 0.144 0.184 0.159 0.208 0.146 0.198 0.147 0.198 0.149 0.198 0.146 0.212 0.149 0.200 0.149 0.200 0.161 0.226 0.172 0.220 0.152 0.237
192 0.186 0.225 0.200 0.248 0.191 0.239 0.192 0.243 0.194 0.241 0.195 0.261 0.193 0.240 0.196 0.245 0.220 0.283 0.219 0.261 0.220 0.282
336 0.237 0.267 0.253 0.289 0.241 0.280 0.247 0.284 0.245 0.282 0.252 0.311 0.241 0.279 0.238 0.277 0.275 0.328 0.280 0.306 0.265 0.319
720 0.307 0.320 0.321 0.338 0.310 0.331 0.318 0.330 0.314 0.334 0.318 0.363 0.324 0.338 0.314 0.334 0.311 0.356 0.365 0.359 0.323 0.362
Avg. 0.219 0.249 0.233 0.271 0.222 0.262 0.226 0.264 0.226 0.264 0.228 0.287 0.226 0.264 0.224 0.264 0.242 0.298 0.259 0.287 0.240 0.300
Electricity 96 0.120 0.214 0.138 0.237 0.132 0.234 0.153 0.256 0.129 0.222 0.135 0.237 0.134 0.228 0.129 0.226 0.159 0.267 0.168 0.272 0.140 0.237
192 0.142 0.237 0.157 0.256 0.148 0.248 0.168 0.269 0.147 0.240 0.160 0.262 0.155 0.248 0.143 0.239 0.168 0.279 0.184 0.289 0.152 0.249
336 0.156 0.252 0.167 0.264 0.165 0.266 0.189 0.291 0.163 0.259 0.182 0.282 0.173 0.268 0.161 0.259 0.196 0.308 0.198 0.300 0.169 0.267
720 0.179 0.278 0.194 0.286 0.197 0.296 0.228 0.320 0.197 0.290 0.246 0.337 0.186 0.282 0.191 0.286 0.203 0.312 0.220 0.320 0.203 0.301
Avg. 0.149 0.245 0.164 0.261 0.161 0.261 0.185 0.284 0.159 0.253 0.181 0.279 0.162 0.256 0.156 0.253 0.182 0.292 0.192 0.295 0.166 0.264
Traffic 96 0.340 0.240 0.363 0.265 0.355 0.261 0.369 0.257 0.360 0.249 0.512 0.282 0.415 0.264 0.368 0.253 0.508 0.301 0.593 0.321 0.410 0.282
192 0.343 0.250 0.385 0.273 0.380 0.271 0.400 0.272 0.379 0.256 0.501 0.273 0.445 0.277 0.379 0.261 0.536 0.315 0.617 0.336 0.423 0.287
336 0.363 0.257 0.396 0.277 0.393 0.281 0.407 0.272 0.392 0.264 0.507 0.279 0.461 0.286 0.397 0.270 0.525 0.310 0.629 0.336 0.436 0.296
720 0.393 0.271 0.445 0.312 0.434 0.300 0.461 0.316 0.432 0.286 0.571 0.301 0.489 0.305 0.440 0.296 0.571 0.323 0.640 0.350 0.466 0.315
Avg. 0.360 0.255 0.397 0.282 0.391 0.278 0.409 0.279 0.391 0.264 0.523 0.284 0.452 0.283 0.396 0.270 0.535 0.312 0.620 0.336 0.434 0.295
Solar 96 0.161 0.224 0.188 0.242 0.165 0.238 0.179 0.232 0.178 0.229 0.166 0.230 0.197 0.241 0.202 0.263 0.188 0.252 0.219 0.314 0.216 0.287
192 0.177 0.237 0.193 0.258 0.184 0.254 0.201 0.259 0.189 0.246 0.186 0.237 0.231 0.264 0.223 0.279 0.215 0.280 0.231 0.322 0.244 0.305
336 0.188 0.244 0.195 0.259 0.191 0.263 0.190 0.256 0.198 0.249 0.203 0.243 0.216 0.272 0.241 0.292 0.222 0.267 0.246 0.337 0.263 0.319
720 0.197 0.252 0.223 0.281 0.199 0.262 0.203 0.261 0.209 0.256 0.210 0.256 0.250 0.281 0.247 0.292 0.226 0.264 0.280 0.363 0.264 0.324
Avg. 0.181 0.239 0.200 0.260 0.185 0.254 0.193 0.252 0.194 0.245 0.191 0.242 0.223 0.264 0.228 0.282 0.213 0.266 0.244 0.334 0.247 0.309
@@ F.2 Ablation Studies
Appendix G Statistical Analysis
@@
Model TimeBridge (Hybrid MAE) TimeBridge (MSE) DeformableTST (2024)
Dataset MSE MAE MSE MAE MSE MAE
ETTm1 0.344 0.379 0.353 0.388 0.348 0.383
ETTm2 0.246 0.310 0.246 0.310 0.257 0.319
ETTh1 0.397 0.424 0.399 0.420 0.404 0.423
ETTh2 0.341 0.382 0.346 0.394 0.328 0.377
Weather 0.219 0.249 0.218 0.250 0.222 0.262
Electricity 0.149 0.245 0.149 0.246 0.161 0.261
Traffic 0.360 0.255 0.360 0.252 0.391 0.278
Solar 0.181 0.239 0.181 0.238 0.185 0.254
Table 15: Average results of different prediction length $O\in\{96,192,336,720\}$ with Hybrid MAE loss and MSE loss.
@@
Table 16: Standard deviation and statistical tests for TimeBridge and second-best method (DeformableTST) on ETT, Weather, Electricity, Traffic and Solar datasets.
@@
Dataset ARR AVol MDD ARR AVol MDD Interval